OpenStudy (anonymous):

Please help! Given an arithmetic sequence with a7=19 and a12=54, find a15.

OpenStudy (anonymous):

75

OpenStudy (anonymous):

Wait! How'd you figure that out?

OpenStudy (anonymous):

U have to form two equations here

OpenStudy (anonymous):

and solve it

OpenStudy (anonymous):

The nth term is a+(n-1)d

OpenStudy (anonymous):

because you have two numbers that are five apart, that means you added the same number to a7 five times to get to a12

OpenStudy (anonymous):

a is the first term

OpenStudy (anonymous):

a12-a7=35 35/5=7 so each term is 7 apart

OpenStudy (anonymous):

then add 7 to a12 three times 54+7+7+7=75